OpenStudy (anonymous):

a sail is in the form of a right triangle that is four times as high as it is wide, the sail is made from 18 square meters of material, what is the height?

OpenStudy (zale101):

Area = 1/2 base x height

OpenStudy (zale101):

\[18 \ =\frac{ 1 }{ 2 }(b)(h)\]

OpenStudy (zale101):

height = 4 * (b) altogether 18 = 1/2 * base * 4 base